write an equation in the slope-intercept form for the line passing through (3, 5) and (8, 15)
Answer by mananth(16949)

x1 y1 x2 y2
3 5 8 15

slope m = (y2-y1)/(x2-x1)
(15-5)/(8-3)
(10/5)
m=2

Plug value of the slope and point (3, 5) in
Y = m x + b
5.00 = 6 + b
b= 5 - 6
b= -1
So the equation will be
Y = 2 x -1
m.ananth@hotmail.ca